# biblio.bib

### v1

@@ Abstract
Complex quantum networks are powerful tools in the modeling of transport phenomena, particularly for biological systems, and enable the study of emergent entanglement structures or topology effects in of many-body quantum systems.
Here, we study the transport properties of a quantum network described by the paradigmatic XXZ Hamiltonian, with non-trivial graph connectivity and topology, and long-range interaction.
Adopting a combination of analytical and numerical methods to analyze the properties of increasingly complex architectures, we find that all-to-all connected regular network preserves over long times the memory of initially injected excitations, tracing it back to the system symmetries and the cooperative shielding. We then develop understanding of the conditions for this property to survive in quantum networks with either power-law node connectivity or complex, small-world type, architectures. Interestingly, we find that memory preserving effects occur also in sparse and more irregular graphs, though to a significantly lower degree.
We discuss the implications of these properties in biology-related problems, such as an application to Weber’s law in neuroscience, and their implementation in specific quantum technologies via biomimicry.
@@ 1 Introduction
Dynamical processes in quantum networks are relevant for the engineering of quantum simulators and quantum computing platforms [Daley2024], for the investigation of biological media under driven-dissipative conditions [plenio2008dephasing, Chisholm_2021, Civolani2024], and to map complex, non-linear phenomena that do not require to be quantum via the novel approach called quantum-like paradigm [YagoMalo2023, khrennikov2010paradigm, basieva2020paradigm]. Complex quantum networks are also becoming powerful tools for extracting relevant information from many-body quantum systems [LCarr], such as emergent entanglement structures, topological instabilities, and self-similarity in quantum states, which may not be apparent by traditional methods [Cirac:2010, Nokkala:2016, VParigi].
@@
The importance of quantum networks, regular or complex, is enhanced by their accurate and precise engineering in current experimental setups, constituting a predominant area of development for quantum simulators [georgescu2014simulation, Altman2024, yagomalo2024review], with widely transversal applications since the first proposals for quantum analog models more than 20 years ago [Daley2024]. Importantly, the current degree of experimental tunability and microscopic understanding of these platforms, both in their coherent and dissipative couplings [daley2014qt], has enabled the quantum simulation of models well beyond condensed matter physics [hofstetter2018condensed] or quantum computing, ranging from fundamental physics and cosmology [Bauer2023], metrology and sensing [pezze2018metro] or – most relevant to this work – quantum thermalization and chaos [eisert2010entropy],[Borgonovi2016] quantum chemistry [cao2019chemistry, Bauer2020] and biology [cao2020biology].
@@
Indeed, the characterization of quantum networks dynamics under open quantum system conditions constitutes a paradigmatic testbed for understanding information processing and transport phenomena in complex media both quantum and classical [biamonte2019complexnetworks], addressing fundamental questions in diverse contexts. These include quantum information theory [swingle2018otoc] with transversal implications for models of quantum gravity [bhattacharyya2022quantum, bentsen2019fast]; material science, in parallel with the experimental and technological development of solid-state nanodevices and cold atom technologies [datta2005transport, Lambert2021]; biological media, where the role of coherent processes in out-of-equilibrium phenomena, though still debated [Fleming2011], has been proven in general complex networks and biological examples [albert2002complexnetworks, ball2011biology, huelga2013vibrations, plenio2008dephasing, vattay2014quantum, biamonte2019complexnetworks]. In fact, in more recent years the study of dynamics in complex networks topologies evolved under paradigmatic Hamiltonians, has raised attention to effectively investigate chemistry and biology applications, by different strategies that can be classified under two main categories. One is a multiscale approach that combines the quantum-chemistry computation of microscopic interaction potentials with the quantum simulation of energy or excitation transport in quantum networks [cao2019chemistry, cao2020biology, ball2011biology]. Another one consists in mapping the complex (classical) network into a quantum network, and exploiting the underlying superposition and entanglement at work to describe the system non-linearities [yagomalo2024review].
@@
The quantum transport properties of a given system are governed by a set of general properties, all of which can be implemented in current quantum technology platforms: (i) the connectivity graph or network topology [Kulvelis2015, mulken2016complexnetworks], where the role of quantum interference produces drastic differences between classical and quantum regimes [biamonte2019complexnetworks]; (ii) the presence of external couplings in the form of driving [Kohler2005, Alamo2024], disorder [Moix_2013, Walschaers2016, Zhou2016] or dissipation [plenio2008dephasing, Wu_2022], and (iii) the coherent Hamiltonian parameters governing the static properties of the model, i.e. not just the transfer rates and interaction strengths, but also the range of such couplings [Kloss2020, Dominguez2021, Defenu2023] with the latter being one of our focuses. Technological development, particularly in ion traps [Bruzewicz2019] and cold gases [Ritsch2013, Adams_2020, Chomaz_2023], has allowed to explore conditions where individual quantum systems are accurately prepared to couple over large distances. The long-range nature of the interactions has been studied in detail and it is known to affect the static and dynamical properties of a quantum model [Defenu2023]. As to the static properties, it has been revealed to enrich the interplay between length scales in the system [Botet:1982, Liu:2019, Defenu:2018], leading to non-additive energy contributions [Campa2009], novel phases of matter, and distinguished changes in known spin [Maghrebi:2017, Giachetti_2021, Defenu2023] or cold-atom systems [Landig:2016, Habibian:2013, Sharma:2022] phase diagrams, than for the case of short-range interactions. As to the dynamical behavior [Campa2009, Dauxois2002], long-range interactions lead to a diversity of interesting phenomena. These concern the occurrence of non-thermal topological steady states [Borgonovi2004, Borgonovi2005], excitation propagation beyond the Lieb-Robinson picture [lieb1972bound, Schachenmayer2013, Eisert2013, Hauke2013], new behavior in the presence of disorder [Chen2023], dissipation [Catalano2023], applications to quantum reservoir computing [sannia2024_skineqnn] or, essential for the current work, trapped dynamics due to interference in what has been denoted as cooperative shielding [santos2016cooperative]. Long-range interactions have revealed to be crucial in preserving over long time both temporal and spatial memory of excitation quenches in a regular, all-to-all connected XXZ open quantum network: this model resulted from the quantum mapping of a classical neuronal complex system and demonstrated [YagoMalo2023] to effectively describe the sense of number [burr2008visual, stoianov2012vision], the very general ability of humans and many animals of non-cognitively counting with constant precision of about 20% up to a few hundred items in a given space or time region [burr2008visual]. Given the paradigmatic nature of the XXZ model  [franchini2017integrable], this time-space memory preserving feature can be predicted to apply to a variety of physical systems, besides being relevant for the development of quantum technology devices.
@@
Understanding the physical mechanism governing this memory-preserving property is therefore of primary importance, exploring the whole variability of transport conditions represented by (i)-(iii) above. To this aim, analytical descriptions are especially valuable in combination with information that can be extracted from quantum simulating the time and spectral signals of the network dynamical observables. In this work, we first provide complete analytical understanding of this phenomenon in regular XXZ quantum networks encompassing (i) with all-to-all connectivity and (ii), after adopting the language of cooperative shielding, where protected subspaces emerge in the evolution due to the system symmetries [santos2016cooperative]. Then, we turn to (iii) and parametrize the extent to which this phenomenon survives in regular quantum networks with power-law varying interactions. We finally revert back to (i) and advance the understanding of this phenomenon in complex XXZ quantum networks, choosing the specific and paradigmatic example of small-world networks, after combining in the semi-analytical outcomes of the so-called cascade model with quantum simulation data. This systematic study sets the stage for a deeper exploration of information and transport processing in paradigmatic XXZ quantum (complex) networks, and the possible presence of universal behavior, revealing some degree of order in disordered structures.
The article is organized as follows. In Section 2, we introduce the XXZ model, analyze its symmetry structure, and exploit the latter to identify the physical mechanism for time-space memory preservation of excitations, in the light of cooperative shielding. The analytics are performed by first studying the all-to-all connected regular XX network and then adding the z-axis anisotropy interaction perturbatively. In Section 3, we explore the same phenomena in all-to-all connected regular networks with power-law decaying interactions and, then, in highly-connected random networks. Finally, in Section 4, we discuss the relevance of our findings for the phenomenological description of complex systems, and the perspective research questions that our work opens up.
@@ 2 Symmetry and dynamics in all-to-all connected quantum networks via cooperative shielding
We begin by introducing the paradigmatic Heisenberg XXZ Hamiltonian we are interested in, with its relevant properties and symmetries. As schematized in Fig. 1$(a)$, the XXZ Hamiltonian is given by:
^$\begin{split}\hat{\mathcal{H}}&=\sum_{(i,j)}\Bigl{[}J_{ij}(\hat{\mathcal{S}}_{i}^{x}\hat{\mathcal{S}}_{j}^{x}+\hat{\mathcal{S}}_{i}^{y}\hat{\mathcal{S}}_{j}^{y})+\Delta_{ij}\hat{\mathcal{S}}_{i}^{z}\hat{\mathcal{S}}_{j}^{z}\Bigr{]}\\ &=\sum_{(i,j)}\Bigl{[}\frac{J_{ij}}{2}(\hat{\mathcal{S}}_{i}^{+}\hat{\mathcal{S}}_{j}^{-}+\hat{\mathcal{S}}_{i}^{-}\hat{\mathcal{S}}_{j}^{+})+\Delta_{ij}\hat{\mathcal{S}}_{i}^{z}\hat{\mathcal{S}}_{j}^{z}\Bigr{]}\,,\end{split}$ (1)
@@
where $\hat{\mathcal{S}}_{i}^{x,y,z}$ represents the local spin magnetization on site $i$ along the given direction and $\hat{\mathcal{S}}_{i}^{\pm}$ represent the corresponding ladder operators for the local spin Hilbert space. The sum runs over all the edges $(i,j)$ of a given graph $G$, representing the interactions between sites. Therefore, for each site the physical neighbors are given by the vertex neighbors on the graph (see Fig. 1$(a)$). Quasi-particles are excited in the system, in the form of spin states (e.g. up, red circles) in a background of opposite spins (e.g. down, blue circles); for simplicity, throughout the paper we will call such quasi-particles excitations, in the sense of a given number of spins up (along the $z$ direction) in a background of spins down as initial state of our system. They couple and travel to connected sites via exchange interaction $J_{ij}$ between spins $i$ and $j$. Sites possessing a link between them are connected by the hopping amplitude $\Delta_{ij}$, which modulates the excitation spatial propagation.
@@
The XXZ model has been studied at length for regular graphs with remarkable differences in local and global observables depending on the degree of coupling [YagoMalo2023, yagomalo2024review, santos2016cooperative, franchini2017integrable]. Here, we focus on the case of highly-connected networks and the important example of all-to-all exchange coupling for the spreading of one single (see Fig. 1$(b)$) and several (see Fig. 1$(d)$) excitations. In the all-to-all connected case, the interference pattern that emerges from the high connectivity leads to substantial differences in the evolution, even at long times, compared to the familiar local models. The completely connected coupling induces high symmetry and leads to a more interesting scenario, since the excitation goes back and forth from the earliest site to all the others in an evenly fashion. From a physical point of view, when the stimulus is located at the initial site $i$, it has equal probability to tunnel to any other node if $\Delta=0$. Each pair of sites $i$ and $j\neq i$ form what can be seen as an effective double-well potential problem, where the part of the excitation shared between that pair tunnels back and forth between the two potential minima. This, regular and symmetric, evolution becomes clear when analyzing the power spectrum of the time signal, Fig. 1$(c)$ and $(e)$, presenting a reduced set of discrete frequencies, as many as the initial excitations are, as we explain in section 2.1. Notice that when considering several excitations, while the time signal profile appears complex (Fig. 1$(d)$), the power spectrum is still well-defined (Fig. 1$(e)$).
Though apparently complicated, these observations can be traced back to an elegant symmetry argument, that we discuss below.
The role of symmetries in highly-connected graphs, Fig. 1$(a)$, for integrable models has paramount importance to understand their quantum transport properties. Some of these dynamical phenomena have been studied in literature in what has been denoted as cooperative shielding [santos2016cooperative]. This phenomenon appears in systems governed by a mean-field Hamiltonian and a long-range perturbation interaction between its constituents. It refers to the existence of protected subspaces within which the evolution remains unaffected for an extended duration due to the appearance of approximate selection rules, with the dynamics heavily depending on the initial state. We exemplify the concept in Fig. 1$(f)$ for a simple situation. Let us consider a Hamiltonian ^=^$\hat{\mathcal{H}}=\hat{\mathcal{H}}_{1}+\hat{\mathcal{H}}_{2}$ for a many-body system, having commuting parts [^$[\hat{\mathcal{H}}_{1},\hat{\mathcal{H}}_{2}]=0$. An initial state $|\psi_{0}\rangle$ belonging to a (degenerate or non-degenerate) eigenspace of ^$\hat{\mathcal{H}}_{2}$ and with corresponding eigenvalue $\lambda_{2}$, evolves in time according to the general Hamiltonian $|\psi(t)\rangle=\text{exp}(-\mathrm{i}\lambda_{2}t)\text{exp}(-\mathrm{i}\hat{\mathcal{H}}_{1}t)|\psi_{0}\rangle$: since ^$\hat{\mathcal{H}}_{2}$ solely induces a global phase without other effects, the dynamics is fully determined by ^$\hat{\mathcal{H}}_{1}$. On the contrary, if the initial state is disseminated across more than one eigenspace of ^$\hat{\mathcal{H}}_{2}$ its dynamics is characterized by $\hat{\mathcal{H}}$. This concept of trapping can be extended to the case of non-commuting Hamiltonian parts in the event of ^$\hat{\mathcal{H}}_{2}$ being composed of solely long-range interactions. If the initial state is initialized in a selected subspace of ^$\hat{\mathcal{H}}_{2}$, long-range interactions do not influence its temporal evolution for a time span that scales with the system size. Instead, the evolution of an initial state with components in more than one subspace, is affected by long-range interactions: the dynamics is not bounded, with components in various subspaces exhibiting fast propagation due to the long-range coupling.
@@
With the cooperative shielding concept at hand, we now analyze in detail the consequences of the Hamiltonian symmetries on the energy levels and their effects on the dynamics. To this aim, we begin with the simplest case of the XX Hamiltonian, i.e. $\Delta_{ij}=0$ in eq. (1), and then we turn it on again as a perturbation.

### v2

# Memory preservation and cooperative shielding in complex quantum networks

@@ Abstract
Complex quantum networks are powerful tools in the modeling of transport phenomena, particularly for biological systems, and enable the study of emergent phenomena in many-body quantum systems. High connectivity and long range interactions induce strong constraints on the system dynamics. Here, we study the transport properties of a quantum network described by the paradigmatic XXZ Hamiltonian, with non-trivial graph connectivity and topology, and long-range interaction. We show how long range interactions induce a memory preserving effects and strongly affect the spreading of the excitations due to cooperative shielding. We describe the memory-preserving effect in all-to-all connected regular networks with distance-independent couplings. Indeed the memory of the number of initially injected excitations is preserved over long times, being encoded in the number of frequencies present in the dynamics. Interestingly, we find that memory-preserving effects occur also in less regular graphs, such as quantum networks with either power-law node connectivity or complex, small-world type, architectures. We discuss the implications of these properties in biology-related problems, such as an application to Weber’s law in neuroscience, and their implementation in specific quantum technologies via biomimicry. We also show how the presence of long range interaction strongly affect the dynamics of the excitations in small-world networks and power law all-to-all coupled networks. Indeed due to cooperative shielding as the connectivity or the range of interaction increase the initial excitation spreads more slowly among the network and becomes strongly dependent on the initial conditions.
@@ 1 Introduction
Dynamical processes in quantum networks are relevant for the engineering of quantum simulators and quantum computing platforms [1], for the investigation of biological media under driven-dissipative conditions [2, 3, 4], and to map complex, non-linear phenomena that do not require to be quantum via the novel approach called quantum-like paradigm [5, 6, 7]. Complex quantum networks are also becoming powerful tools for extracting relevant information from many-body quantum systems [8], such as emergent entanglement structures, topological instabilities, and self-similarity in quantum states, which may not be apparent by traditional methods [9, 10, 11].
@@
The importance of quantum networks, regular or complex, is enhanced by their accurate and precise engineering in current experimental setups, constituting a predominant area of development for quantum simulators [12, 13, 14], with widely transversal applications since the first proposals for quantum analog models more than 20 years ago [1]. Importantly, the current degree of experimental tunability and microscopic understanding of these platforms, both in their coherent and dissipative couplings [15], has enabled the quantum simulation of models well beyond condensed matter physics [16] or quantum computing, ranging from fundamental physics and cosmology [17], metrology and sensing [18] or – most relevant to this work – quantum thermalization and chaos [19],[20] quantum chemistry [21, 22] and biology [23].
@@
Indeed, the characterization of quantum networks dynamics under open quantum system conditions constitutes a paradigmatic testbed for understanding information processing and transport phenomena in complex media both quantum and classical [24], addressing fundamental questions in diverse contexts. These include quantum information theory [25] with transversal implications for models of quantum gravity [26, 27]; material science, in parallel with the experimental and technological development of solid-state nanodevices and cold atom technologies [28, 29]; biological media, where the role of coherent processes in out-of-equilibrium phenomena, though still debated [30], has been proven in general complex networks and biological examples [31, 32, 33, 2, 34, 24]. In fact, in more recent years the study of dynamics in complex networks topologies evolving under paradigmatic Hamiltonians, has raised attention to effectively investigate chemistry and biology applications, by different strategies that can be classified under two main categories. One is a multiscale approach that combines the quantum-chemistry computation of microscopic interaction potentials with the quantum simulation of energy or excitation transport in quantum networks [21, 23, 32]. Another one consists in mapping the complex (classical) network into a quantum network, and exploiting the underlying superposition and entanglement at work to describe the system nonlinearities [14].
@@
The quantum transport properties of a given system are governed by a set of general properties, all of which can be implemented in current quantum technology platforms: (i) the connectivity graph or network topology [35, 36], where the role of quantum interference produces drastic differences between classical and quantum regimes [24]; (ii) the presence of external couplings in the form of driving [37, 38], disorder [39, 40, 41] or dissipation [2, 42], and (iii) the coherent Hamiltonian parameters governing the static properties of the model, i.e. not just the transfer rates and interaction strengths, but also the range of such couplings [43, 44, 45] with the latter being one of our focuses. Technological development, particularly in ion traps [46] and cold gases [47, 48, 49], has allowed to explore conditions where individual quantum systems are accurately prepared to couple over large distances. The long-range nature of the interactions has been studied in detail and it is known to affect the static and dynamical properties of a quantum model [45]. As to the static properties, it has been revealed to enrich the interplay between length scales in the system [50, 51, 52], leading to non-additive energy contributions [53], novel phases of matter, and distinguished changes in spin systems [54, 55, 45] or cold-atom systems [56, 57, 58] phase diagrams, than for the case of short-range interactions. As to the dynamical behavior [53, 59], long-range interactions lead to a diversity of interesting phenomena. These concern the occurrence of non-thermal topological steady states [60, 61], excitation propagation beyond the Lieb-Robinson picture [62, 63, 64, 65], new behavior in the presence of disorder [66, 67], dissipation [68], applications to quantum reservoir computing [69] or, essential for the current work, trapped dynamics due to interference in what has been denoted as cooperative shielding [70, 71].
This phenomenon, which will be relevant in the following, appears in systems with long-range interaction between its constituents. It refers to the existence of protected subspaces within which the evolution remains unaffected for a finite time due to the appearance of approximate selection rules, with the dynamics heavily depending on the initial state. To better understand what is meant by shielding, we exemplify the concept for a simple situation. Let us consider a Hamiltonian $\hat{\mathcal{H}}=\hat{\mathcal{H}}_{1}+\hat{\mathcal{H}}_{2}$ for a many-body system, having commuting parts $\hat{\mathcal{H}}_{1},\hat{\mathcal{H}}_{2}]=0$. An initial state $|\psi_{0}\rangle$ belonging to a (degenerate or non-degenerate) eigenspace of $\hat{\mathcal{H}}_{2}$ and with corresponding eigenvalue $\lambda_{2}$, evolves in time according to the general Hamiltonian $|\psi(t)\rangle=\text{exp}(-\mathrm{i}\lambda_{2}t)\text{exp}(-\mathrm{i}\hat{\mathcal{H}}_{1}t)|\psi_{0}\rangle$: since $\hat{\mathcal{H}}_{2}$ solely induces a global phase without other effects, the dynamics is fully determined by $\hat{\mathcal{H}}_{1}$. So in this case we can say that the dynamics is shielded from the effects of $\hat{\mathcal{H}}_{2}$. On the contrary, if the initial state is disseminated across more than one eigenspace of $\hat{\mathcal{H}}_{2}$ its dynamics is characterized by $\hat{\mathcal{H}}$.
The concept of shielding can be extended to the case of non-commuting Hamiltonian parts in the event of $\hat{\mathcal{H}}_{2}$ being composed of solely long-range interactions. If the initial state is initialized in a selected subspace of $\hat{\mathcal{H}}_{2}$, long-range interactions do not influence its temporal evolution for a time span that scales with the system size. Instead, the evolution of an initial state with components in more than one subspace, is affected by long-range interactions: the dynamics is not bounded, with components in various subspaces exhibiting fast propagation due to the long-range coupling. Cooperative shielding can induce an effective localization of the initial excitation in long-range interacting system. This effect is counter-intuitive since one could expect that the larger the range of the interaction is, the easier the initial excitation will spread away from its initial position. Here we show that this effect has a considerable impact on the dynamics of highly connected networks.
Long-range interactions have revealed to be crucial in preserving over long time both temporal and spatial memory of excitation quenches in a regular, all-to-all connected XXZ open quantum network: this model resulted from the quantum mapping of a classical neuronal complex system and demonstrated [5] to effectively describe the sense of number [72, 73], the very general ability of humans and many animals of non-cognitively counting with constant precision of about 20% up to a few hundred items in a given space or time region [72]. Given the paradigmatic nature of the XXZ model  [74], this time-space memory preserving feature can be predicted to apply to a variety of physical systems, besides being relevant for the development of quantum technology devices.
@@
Understanding the physical mechanism governing this memory-preserving property is therefore of primary importance, especially when considering the whole variability of transport conditions represented by (i)-(iii) above. To this aim, analytical descriptions are particularly valuable in combination with information that can be extracted from quantum simulating the time and spectral signals of the network dynamical observables. In this work, we first provide complete analytical understanding of this phenomenon in regular, all-to-all connected XXZ quantum networks, in whih symmetries play a pivotal role. Then, we turn to the aforementioned points (i) and (iii) to parametrize the extent to which this phenomenon survives in more general cases: we first devote the analysis to complex quantum networks described by the XXZ Hamiltonian having couplings based on a small-world architecture; secondly, we move back to completely connected quantum networks with power-law varying interactions. In both cases, cooperative shielding and long-range interactions play a fundamental role in the memory preservation effect through non-trivial modalities. This systematic study sets the stage for a deeper exploration of information and transport processing in paradigmatic XXZ quantum (complex) networks, and the possible presence of universal behavior, revealing some degree of order in disordered structures. The article is organized as follows. In Section 2, we introduce the XXZ model, analyze its symmetry structure, and exploit the latter to identify the physical mechanism for time-space memory preservation of excitations, in the light of cooperative shielding. The analytics are performed by first studying the all-to-all connected regular XX network and then adding the $z$-axis anisotropy interaction perturbatively. In Section 3, we explore the same phenomena in highly-connected random networks and, then, in all-to-all connected regular networks with power-law decaying interactions. Finally, in Section 4, we discuss the relevance of our findings for the phenomenological description of complex systems, and the perspective research questions that our work opens up.
@@ 2 Symmetry and dynamics in all-to-all connected quantum networks via cooperative shielding
We begin by introducing the paradigmatic Heisenberg XXZ Hamiltonian we are interested in, with its relevant properties and symmetries; throughout the paper we will work in natural units, i.e. $\hbar=1$. As schematized in Fig. 1(a)(a), the XXZ Hamiltonian is given by:
[…]
@@
where $\hat{\mathcal{S}}_{i}^{x,y,z}$ represents the local spin-1/2 magnetization on site $i$ along the given direction and $\hat{\mathcal{S}}_{i}^{\pm}$ represent the corresponding ladder operators for the local spin-1/2 Hilbert space. The sum runs over all the edges $(i,j)$ of a given graph $G$, representing the interactions between sites. Therefore, for each site the physical neighbors are given by the vertex neighbors on the graph (see Fig. 1(a)(a)). Quasi-particles are excited in the system, in the form of spin states (e.g. up, red circles) in a background of opposite spins (e.g. down, blue circles); for simplicity, throughout the paper we will call such quasi-particles excitations. The number of excitations indicate the number of spins up (along the $z$ direction) in a background of spins down. Excitations can hop from site to site via exchange interaction $J_{ij}$ between spins $i$ and $j$. The last term $\Delta_{ij}$ represents an additional interaction between the $z$-components of the spins.
The XXZ model has been studied at length for regular graphs with remarkable differences in local and global observables depending on the degree of coupling [5, 14, 70, 74]. Here, we focus on the case of highly-connected networks. We also focus on the important example of all-to-all exchange coupling (where all the $J_{ij}$ are the same independently of the distance) for different initial conditions, see Fig. 1(b)(b) Fig. 1(d)(d). The initial condition is taken to be an eigenstate of the total $z$-magnetization operator $\hat{\mathcal{M}}_{z}=\sum_{i}\hat{\mathcal{S}}_{i}^{z}$, which assume the general form $|\boldsymbol{\sigma}_{q}\rangle=|\sigma_{1}^{q}\dots\sigma_{L}^{q}\rangle$, with each $\sigma$ denoting either spin-up, $\uparrow$, or spin-down, $\downarrow$; this string will be dubbed as Pauli string, while the basis set $\{|\boldsymbol{\sigma}_{q}\rangle,q=1,\dots,2^{L}\}$ will be called computational basis.
@@
In the all-to-all connected case, the interference pattern that emerges from the high connectivity leads to substantial differences in the evolution, even at long times, compared to the familiar local models. From a physical point of view, when the excitation is located at the initial site $i$, it has equal probability to tunnel to any other node if $\Delta=0$. This leads to a regular and symmetric evolution as it is seen in the power spectrum of the time signal, Fig. 1(c)(c) and $(e)$. Notably, the power spectrum shows a number of discrete frequencies equal to the number of initial excitations, see section 2.1. Notice that when considering several excitations, while the time signal profile appears complex (Fig. 1(d)(d)), the power spectrum is still well-defined (Fig. 1(e)(e)).
@@
The role of symmetries in highly-connected graphs, Fig. 1(a)(a), for integrable models has paramount importance to understand their quantum transport properties. We now analyze in detail the consequences of the Hamiltonian symmetries on the energy levels and their effects on the dynamics. To this aim, we begin with the simplest case of the XX Hamiltonian, i.e. $\Delta_{ij}=0$ in eq. (1), and then we turn it on again as a perturbation.